OpenStudy (anonymous):

Let f(x) = x + 7 and g(x) = x – 4. Find f(x) • g(x).

OpenStudy (anonymous):

(x+7)(x-4)

OpenStudy (anonymous):

f(x).g(x)=(x+7)(x-4) =x(x-4)+7(x-4) hope you can solve it further........

OpenStudy (jiteshmeghwal9):

\[f(x).g(x)\]\[=(x+7)(x-4)\]\[=x(x-4)+7(x-4)\]\[=x^2-4x+7x-28\]\[=x^2+3x-28\]